# RFOX (Rotated-Field Oscillatory eXchange) quantum algorithm: Towards Parameter-Free Quantum Optimizers

### 02 Apr 2026 v1

@@ Abstract
We introduce RFOX (Rotated-Field Oscillatory eXchange), a parameter-free quantum algorithm for combinatorial optimization. RFOX combines an almost constant non-stoquastic $XX$ catalyst with a weak harmonic $ZX$ counter-diabatic term. Using the Floquet-Magnus expansion, we derive a closed-form effective Hamiltonian whose first-order term retains the full $XX$ driver, while the leading correction consists of a single qubit $Y$ field at high drive frequency. This structure ensures that the instantaneous spectral gap remains essentially flat, independent of both the interpolation parameter and the disorder strength, modulated only by a $\delta$ parameter. This behavior stands in stark contrast to the unpredictable gap reductions, or even collapses, exhibited by the XX (stoquastic), $XX$, and $X+sXX$ (non-stoquastic) driver schedules. Extensive noiseless simulations on random-field Ising model (RFIM) instances with 7, 9, and 12 qubits, across three magnetic-field ranges, validate these spectral predictions: RFOX attains near-optimal, and in some cases exact, ground states using up to an order of magnitude fewer Trotter slices. Its performance advantage grows with increasing disorder, as conventional methods slow down near vanishing gaps, whereas RFOX maintains a constant runtime scaling of $T\propto\Delta_{\min}^{-2}$. Hardware experiments on IBM Quantum processors (Eagle r3 and Heron r1, with 12, 15, and 20 physical qubits) reproduce similar performance rankings. These results suggest that fixed-gap, non-stoquastic drivers augmented with analytically derived counter-diabatic terms offer a promising, scalable, and tuning-free route toward quantum optimizers for combinatorial optimization problems.
@@ I Introduction
In disordered systems such as the RFIM, where frustrated couplings and random longitudinal fields hinder conventional quantum annealing, introducing a transverse interaction channel $X\otimes X$ (XX) can accelerate convergence and suppress localization-induced gap closures, outperforming pure transverse field schedules Suzuki et al. (2007); Zhou and others (2022); Yang and others (2024). A key design element of RFOX is the combination of a nearly constant non‐stoquastic XX driver with a counter-diabatic term of time oscillation $Z\otimes X$ (ZX). The XX interaction alleviates the stoquastic bottleneck by coupling computational basis states differing by multiple spin flips, thus widening the spectral gap Lykiardopoulou et al. (2021); Albash (2019); Choi (2021). However, a large gap alone does not guarantee adiabatic evolution: when problem fields vary rapidly, residual diabatic transitions can accumulate. The harmonic ZX term serves as the leading-order component of the adiabatic gauge potential, modulated by a phase-shifted sinusoidal envelope that supplies precisely the time derivative required to cancel non-adiabatic couplings, without adding any static cost term that could close the gap. As a result, the combined XX+ZX driver preserves a nearly constant minimum gap while suppressing diabatic leakage, enabling the system to reach approximate, and in some cases exact, ground states at fixed, shallow depth and without classically tuned parameters.
@@ II.1 Magnetic field encoding using phase gates
  • Since $B_{ZX}(k)$ is antisymmetric with zero mean, it serves as a first-order counter-diabatic (CD) correction, canceling diabatic transitions induced by $H_{XX}$ to leading order without altering the instantaneous spectrum Cárdenas-López and Chen (2022); Claeys et al. (2019).
@@
  • The factor of $Z_{u}$ ties each kick’s amplitude to the current spin configuration, biasing collective spin flips toward low‑energy valleys in the RFIM landscape.
Because $\sum_{k=0}^{p-1}B_{ZX}(k)=0$, the total number of slices remains $p$, and the circuit depth scales as $2p\lvert E\rvert$, independent of any adiabatic‐schedule parameter.
II.3 Counter‑diabatic driving and the adiabatic gauge potential
@@
producing off-diagonal couplings that exactly cancel the diabatic excitations. For spin systems, $A_{\lambda}$ is an imaginary Hermitian operator and typically contains an odd number of Pauli‑$Y$ matrices Sels and Polkovnikov (2017).
RFOX implements this CD mechanism via its time‑dependent ZX operator. Specifically:
  • The second-order Magnus term ($i\,\frac{\delta^{2}}{\omega}\sum_{u}Y_{u}$) emerges naturally from the fast, harmonic ZX drive and matches the local AGP (AGP consistency) for an Ising-type system driven by a transverse field. The single $Y$ Pauli terms anticommutes with both the transverse-field term in $H_{B}$ and the XX backbone, providing the minimal transition set needed to cancel diabatic excitations.
  • Instead of adding $\sum_{u}Y_{u}$ explicitly, RFOX synthesises the CD term: the $\sin(\omega t)ZX$ kick, combined with the static XX interaction, generates $i\frac{\delta^{2}}{\omega}\sum_{u}Y_{u}$ via the commutator $[XX,ZX]$ in the Magnus expansion. This produces the CD correction without enlarging the native control set.
  • The Floquet‑engineered $Y$ term prevents gap closures along the evolution path by converting would-be level crossings into avoided crossings. Numerical simulations and prior CD-engineering studies confirm that this mechanism maintains a near-constant minimum gap and high ground-state fidelity at finite runtimes Cárdenas-López and Chen (2022); Claeys et al. (2019); Sels and Polkovnikov (2017).
@@ III.3 Floquet-Magnus effective Hamiltonian
$\Omega_{1}=H_{B}+H_{XX}\ .$ (22)
For the second Magnus term we evaluate the commutator of $H_{XX}$ and $H_{ZX}$ given as:
$\left[H_{XX},H_{ZX}\right]=\sum_{(u,v)\in E}\left[X_{u}X_{v},Z_{u}X_{v}\right]=\sum_{(u,v)\in E}X_{u}X_{v}(Z_{u}X_{v})-Z_{u}X_{v}(X_{u}X_{v})=\sum_{(u,v)\in E}-2iY_{u}\ ,$ (23)
then we take the time-dependent parts
$\left[H_{osc}(t_{1}),H_{osc}(t_{2})\right]=\delta^{2}\sin\left[\omega(t_{1}-t_{2})\right]\left[H_{XX},H_{ZX}\right]=2i\delta^{2}\sin\left[\omega(t_{1}-t_{2})\right]Y_{u}\ .$ (24)
We omit the $\sum_{(u,v)\in E}$ part to simply the expression, but the term $Y_{u}$ maintains the summation over all edges. The second Magnus term is given by
$\Omega_{2}=\frac{1}{2}\int_{0}^{T}dt_{1}\int_{0}^{t_{1}}dt_{2}\ \left(-2i\delta^{2}\sin\bigl[\omega(t_{1}-t_{2})\bigr]\right)Y_{u}=-i\delta^{2}\Bigl[\frac{T}{\omega}\Bigr]\sum_{(u,v)\in E}Y_{u}.$ (25)
we have the effective Hamiltonian to $\mathcal{O}(\delta^{2})$:
$H_{\rm eff}=H_{B}+\sum_{(u,v)\in E}X_{u}X_{v}-i\frac{\delta^{2}}{\omega}\sum_{(u,v)\in E}Y_{u}+\mathcal{O}(\delta^{3}/w^{2})\ .$ (26)
RFOX maintains a robust non-stoquastic XX backbone that prevents gap closures, while a high-frequency ZX kick serves as a CD pulse Schindler and Bukov (2024); Claeys et al. (2019). In this way, the phase information encoding the RFIM cost accumulates without collapsing the spectral gap, and the final projective measurement produces a bit-string distribution sharply peaked at the ground state.
@@ VI Conclusions
Our numerical study demonstrates that the RFOX schedule—an always-on, non-stoquastic XX catalyst augmented by a weak, harmonic ZX counter-diabatic kick—consistently outperforms both the X-only and XX drivers, and the commonly used X+sXX catalyst. Across all tested random-field Ising model instances (7, 9, and 12 qubits; three magnetic-field ranges; Erdös–Rényi and Watts–Strogatz graphs), RFOX achieves the lowest cost average, the smallest excess energy above the exact optimum, and the shortest average Hamming distance to the ground-state bitstring. These performance gains directly reflect the Floquet-Magnus analysis: the first-order term preserves the full strength of the non-stoquastic XX driver, providing a robust gap floor, while the second-order term introduces only a perturbative $Y$-field micromotion counter-diabatic correction that refines, rather than narrows, the spectrum. Consequently, the minimum gap remains nearly constant throughout the evolution, whereas comparison schedules exhibit slow starts, gap shrinkage, or even pronounced gap collapse. Since adiabatic runtime scales as $\Delta_{\min}^{-2}$, this nearly flat-gap profile enables RFOX to approach the ground state in far fewer Trotter slices, and thus in shallower circuits, than its competitors.
Hardware experiments on Eagle-class (ibm_brisbane and ibm_sherbrooke) and Heron-class (ibm_torino) backends confirm that RFOX’s advantage persists under realistic noise. Although decoherence increases errors across all metrics, RFOX consistently outperforms baseline schedules. This robustness is partly attributable to amplitude-damping noise, which shifts probability weight toward low-energy configurations, a shift that RFOX’s large minimum gap helps preserve. Importantly, RFOX delivers these improvements with shallower circuits as those of the comparator methods and requires no classical parameter optimization.

### 06 May 2026 v2

@@ Abstract
We introduce RFOX (Rotated-Field Oscillatory eXchange), a parameter-free quantum algorithm for combinatorial optimization. RFOX combines an almost constant non-stoquastic $XX$ catalyst with a weak harmonic $ZX$ counter-diabatic term. Using the Floquet-Magnus expansion, we derive a closed-form effective Hamiltonian whose first-order term retains the full $XX$ driver, while the leading correction consists of a local single-qubit $Y$ field and poly-local 3-body topological interactions driven by the graph connectivity at high drive frequency. This structure ensures that the instantaneous spectral gap remains essentially flat, independent of both the interpolation parameter and the disorder strength, modulated only by a $\delta$ parameter. This behavior stands in stark contrast to the unpredictable gap reductions, or even collapses, exhibited by the XX (stoquastic), $XX$, and $X+sXX$ (non-stoquastic) driver schedules. Extensive noiseless simulations on random-field Ising model (RFIM) instances with 7, 9, and 12 qubits, across three magnetic-field ranges, validate these spectral predictions: RFOX attains near-optimal, and in some cases exact, ground states using up to an order of magnitude fewer Trotter slices. Its performance advantage grows with increasing disorder, as conventional methods slow down near vanishing gaps, whereas RFOX maintains a constant runtime scaling of $T\propto\Delta_{\min}^{-2}$. Hardware experiments on IBM Quantum processors (Eagle r3 and Heron r1, with 12, 15, and 20 physical qubits) reproduce similar performance rankings. These results suggest that fixed-gap, non-stoquastic drivers augmented with analytically derived counter-diabatic terms offer a promising, scalable, and tuning-free route toward quantum optimizers for combinatorial optimization problems.
@@ I Introduction
In disordered systems such as the RFIM, where frustrated couplings and random longitudinal fields hinder conventional quantum annealing, introducing a transverse interaction channel $X\otimes X$ (XX) can accelerate convergence and suppress localization-induced gap closures, outperforming pure transverse field schedules Suzuki et al. (2007); Zhou and others (2022); Yang and others (2024). A key design element of RFOX is the combination of a nearly constant non‐stoquastic XX driver with a counter-diabatic term of time oscillation $Z\otimes X$ (ZX). The XX interaction alleviates the stoquastic bottleneck by coupling computational basis states differing by multiple spin flips, thus widening the spectral gap Lykiardopoulou et al. (2021); Albash (2019); Choi (2021). However, a large gap alone does not guarantee adiabatic evolution: when problem fields vary rapidly, residual diabatic transitions can accumulate. The harmonic ZX term serves as the leading-order component of the adiabatic gauge potential, modulated by a phase-shifted sinusoidal envelope. As derived via the Floquet-Magnus expansion, this supplies precisely the time derivative required to cancel non-adiabatic couplings by generating both local spin rotations and poly-local topological interactions dictated by the graph connectivity, without adding any static cost term that could close the gap. As a result, the combined XX+ZX driver preserves a nearly constant minimum gap while suppressing diabatic leakage, enabling the system to reach approximate, and in some cases exact, ground states at fixed, shallow depth and without classically tuned parameters.
@@ II.1 Magnetic field encoding using phase gates
  • Since $B_{ZX}(k)$ is antisymmetric with zero mean, it serves as a first-order counter-diabatic (CD) correction. Through its high-frequency commutation with the $XX$ backbone, it cancels diabatic transitions by effectively generating both local spin rotations and poly-local topological interactions without altering the instantaneous static spectrum Cárdenas-López and Chen (2022); Claeys et al. (2019).
@@
  • The factor of $Z_{u}$ ties each kick’s amplitude to the current spin configuration. Because the $Z$ operator does not commute with $X$ operators on incident edges, this structure intrinsically binds collective spin flips to the graph connectivity, generating the 3-body topological correlations that guide the system toward low-energy valleys in the RFIM landscape.
Because $\sum_{k=0}^{p-1}B_{ZX}(k)=0$, the total number of slices remains $p$, and the circuit depth scales as $2p\lvert E\rvert$, independent of any adiabatic-schedule parameter.
II.3 Counter-diabatic driving and the adiabatic gauge potential
@@
producing off-diagonal couplings that exactly cancel the diabatic excitations. For spin systems, $A_{\lambda}$ is an imaginary Hermitian operator and typically contains an odd number of Pauli-$Y$ matrices Sels and Polkovnikov (2017).
RFOX implements this CD mechanism via its time-dependent ZX operator. Specifically:
  • The second-order Magnus term $\left(-\frac{\delta^{2}}{\omega}\sum_{(u,v)\in\vec{E}}\left(Y_{u}+\sum_{w\in N(u)\setminus\{v\}}Y_{u}X_{v}X_{w}\right)\right)$ emerges naturally from the fast, harmonic ZX drive and acts as an approximate local AGP for the Ising-type system. These real and Hermitian terms anticommute with both the transverse-field term in $H_{B}$ and the XX backbone, providing the necessary transition set to cancel diabatic excitations.
  • Instead of adding these terms explicitly, RFOX synthesises the CD correction: the $\sin(\omega t)ZX$ kick, combined with the static XX interaction, generates both the local $Y_{u}$ rotations and the poly-local 3-body topological interactions via the commutator $[XX,ZX]$ in the Magnus expansion. This produces the CD correction without enlarging the native control set.
  • The Floquet-engineered effective fields prevent gap closures along the evolution path by converting would-be level crossings into avoided crossings. Numerical simulations and prior CD-engineering studies confirm that this mechanism maintains a near-constant minimum gap and high ground-state fidelity at finite runtimes Cárdenas-López and Chen (2022); Claeys et al. (2019); Sels and Polkovnikov (2017).
@@ III.3 Floquet-Magnus effective Hamiltonian
$\Omega_{1}=(H_{B}+H_{XX})T\ .$ (22)
For the second Magnus term, we evaluate the commutator of $H_{XX}$ and $H_{ZX}$. Because $H_{ZX}$ applies $Z_{u}$ and $X_{v}$ asymmetrically, we denote the directed edge set as $\vec{E}$, where $u$ is the target of the $Z$ operator. The operator $Z_{u}$ does not commute with any $X_{u}$ acting on the same node, meaning we must sum over all adjacent neighbors $w\in N(u)$:
$H_{XX},H_{ZX}]=\sum_{(u,v)\in\vec{E}}\left(\sum_{w\in N(u)}[X_{u}X_{w},Z_{u}X_{v}]\right)$ (23)
Evaluating this yields a local term when $w=v$, and poly-local 3-body interactions when $w\neq v$:
$H_{XX},H_{ZX}]=-2i\sum_{(u,v)\in\vec{E}}Y_{u}-2i\sum_{(u,v)\in\vec{E}}\left(\sum_{w\in N(u)\setminus\{v\}}Y_{u}X_{v}X_{w}\right)$ (24)
Taking the time-dependent parts of the Magnus expansion and integrating over the time-ordered domain gives the second term $\Omega_{2}$:
$\Omega_{2}=-\frac{i}{2}\int_{0}^{T}dt_{1}\int_{0}^{t_{1}}dt_{2}\left(-2i\delta^{2}\sin[\omega(t_{1}-t_{2})]\right)\frac{[H_{XX},H_{ZX}]}{-2i}=-\delta^{2}\left[\frac{T}{\omega}\right]\sum_{(u,v)\in\vec{E}}\left(Y_{u}+\sum_{w\in N(u)\setminus\{v\}}Y_{u}X_{v}X_{w}\right)$ (25)
Dividing $\Omega_{1}+\Omega_{2}$ by the period $T$, we obtain the effective Hamiltonian to $\mathcal{O}(\delta^{2})$:
$H_{eff}=H_{B}+\sum_{e\in E}X_{e_{a}}X_{e_{b}}-\frac{\delta^{2}}{\omega}\sum_{(u,v)\in\vec{E}}Y_{u}-\frac{\delta^{2}}{\omega}\sum_{(u,v)\in\vec{E}}\left(\sum_{w\in N(u)\setminus\{v\}}Y_{u}X_{v}X_{w}\right)+\mathcal{O}(\delta^{3}/\omega^{2})$ (26)
To simplify the analysis, isolating the term $\frac{\delta^{2}}{\omega}\sum Y_{u}$ can be treated as a truncated mean-field approximation that captures the primary counter-diabatic (CD) Schindler and Bukov (2024); Claeys et al. (2019) local spin rotations. We propose that the residual 3-body terms ($Y_{u}X_{v}X_{w}$) are not negligible noise, but rather represent vital topological correlations introduced by the graph’s connectivity. Because RFOX naturally implements the exact discrete gates $U(\theta,\phi)$ on hardware, these 3-body couplings are inherently executed. This topological awareness likely contributes to the algorithm’s performance advantage: the CD pulses assist in breaking frustrated local configurations dictated specifically by the graph structure, mitigating stochastic errors and maintaining a robust non-stoquastic XX backbone that prevents gap closures.
@@ VI Conclusions
Our numerical study demonstrates that the RFOX schedule-an always-on, non-stoquastic $XX$ catalyst augmented by a weak, harmonic $ZX$ counter-diabatic kick, consistently outperforms both the XX-only and $XX$ drivers, and the commonly used $X+sXX$ catalyst. Across all tested random-field Ising model instances (7, 9, and 12 qubits; three magnetic-field ranges; Erdös–Rényi and Watts–Strogatz graphs), RFOX achieves the lowest cost average, the smallest excess energy above the exact optimum, and the shortest average Hamming distance to the ground-state bitstring. These performance gains directly reflect the Floquet-Magnus analysis: the first-order term preserves the full strength of the non-stoquastic $XX$ driver, providing a robust gap floor, while the second-order term introduces a perturbative counter-diabatic correction comprising both a local $Y$-field and poly-local 3-body topological interactions. This structure refines, rather than narrows, the spectrum and intrinsically maps the graph connectivity into the evolution. Consequently, the minimum gap remains nearly constant throughout the evolution, whereas comparison schedules exhibit slow starts, gap shrinkage, or even pronounced gap collapse. Since adiabatic runtime scales as $\Delta_{\min}^{-2}$, this nearly flat-gap profile enables RFOX to approach the ground state in far fewer Trotter slices, and thus in shallower circuits, than its competitors.
Hardware experiments on Eagle-class (ibm_brisbane and ibm_sherbrooke) and Heron-class (ibm_torino) backends confirm that RFOX’s advantage persists under realistic noise. Although decoherence increases errors across all metrics, RFOX consistently outperforms baseline schedules. This robustness is partly attributable to amplitude-damping noise, which shifts probability weight toward low-energy configurations, a shift that RFOX’s large minimum gap helps preserve. Importantly, RFOX delivers these improvements with shallower circuits than those of the comparator methods and requires no classical parameter optimization.